# Safe Reinforcement Learning with Free-form Natural Language Constraints and Pre-Trained Language Models

### v1

@@ Abstract.
Reinforcement Learning, Natural Language Constraints, Language Models
@@ 1. Introduction
Recent years has seen great improvement in reinforcement learning (RL) algorithms on multiple domains such as robotic control Kaufmann et al. (2023), resource allocation Zheng et al. (2022) and achieving expert level in games Vinyals et al. (2019); Perolat et al. (2022). In real-world settings such as autonomous driving Gu et al. (2022), RL agents should not be given complete freedom for safety, fairness or other concerns and must adhere to specific constraints provided by humans. To this end, safe RL follows the constrained Markov Decision Process (CMDP) Altman (1999) framework, where agents must satisfy the constraints given by auxiliary costs when optimizing their policies to maximize the objective function.
@@
To learn a safe policy, safe RL algorithms Achiam et al. (2017); Ray et al. (2019); Yang et al. (2020, 2022) utilize ground-truth cost value given by a well-defined cost function in CMDP. Because the conditions for violating constraints are different, cost functions must be defined for all constraints case by case and require task-specific domain knowledge, which can be very expensive and inefficient. For example, in a navigation task, the agent should get close to objects that may cause damage to it. To define such a cost function, one must be aware of all the harmful objects in the environment, otherwise the agent may be harmed by some unexpected hazards. This severely limits safe RL’s application to more domains as transforming constraints into cost functions is expensive and requires specific domain knowledge.
@@
One intuitive and easily-accessible way to constrain RL agents is to provide them constraints or instructions via human language Kaplan et al. (2017); Prakash et al. (2020); Yang et al. (2021). By providing natural language constraints, potential end users can easily regularize and interact with agents. However, without a ground-truth cost function, it is difficult to determine whether natural language constraints are violated. To avoid using ground-truth costs, previous method Yang et al. (2021) learns to identify entities relevant to the natural language constraints in the environment so that agents can predict whether the constraints are violated. However, it requires additional computation to model the entities and may result in inaccurate predictions in complex environments. And since the model is trained in a supervised fashion, it is incapable of handling novel and free-form natural language from humans.
To deal with the problem of constraint violation prediction in safe RL with natural language constraints, we propose to adopt pre-trained language models (LM) Devlin et al. (2018); Brown et al. (2020) to comprehend and encode the constraints to constrain the policy and predict costs for policy learning.
@@
Specifically, a decoder-based LM (such as GPT Brown et al. (2020)) is adopted to eliminate ambiguity and redundancy and extract semantic meaning from the constraints. As decoder LMs are pre-trained on a very large corpus of data and align with human values, they are able to generate high-quality constraints with condensed semantic meaning if we prompt them properly. To condition the policy which is parameterized by neural networks with natural language constraints, an encoder-based LM (such as BERT Devlin et al. (2018)) encodes constraints into embeddings according to their semantic meaning. We also use the encoder LM to encode the text-based observations. By computing the semantic similarities between constraint embeddings and description embeddings, we can determine whether the constraint is violated. To distinct constraint embeddings from each other, we adopt a contrastive loss to fine-tune the encoder LM so that constraints have similar embeddings only when they have similar semantic meaning.
@@
    We introduce pre-trained LMs to safe RL with natural language constraints to replace the recurrent neural network in previous works, which is only able to handle structured or less free-form natural language constraints.
@@ 4.1. Cost Prediction Module
We give the structure of the cost prediction module in Fig. 1.
In general, to predict whether the constraint is violated, we adopt a decoder LM (GPT) to eliminate ambiguity and condense semantic meaning in free-form natural language constraint $c$ and an encoder LM (BERT) to extract semantic information from the condensed natural language constraint $\overline{c}$ and transform the constraints into embedding $h_{c}$. Condensed constraint $\overline{c}$ aligns with the original constraint $c$ on which entity or behaviour the agents are prohibited from. This is necessary as free-form constraint $c$ from humans may be verbose or semantically vague, which will severely affect the cost prediction result. Since decoder LM GPT aligns with human values Ziegler et al. (2019), it can eliminate ambiguity within the human natural language constraints and better summarize the constraints. We empirically show in the experiment section that without the decoder LM, the performance of the cost prediction module will drop significantly. The prompt we use for the decoder LM is provided in the supplementary material.
@@ 4.2. Policy Training
Algorithm 1 PPO with LM-based Cost Prediction (PPO_CP)
@@
$L=L^{CLIP}+\alpha L^{CLIP}_{C}-c_{1}L^{VF}-c_{2}L^{CVF},$ (8)
@@ 5. Experiment
Figure 3. Experiment results on Hazard-World-Grid and SafetyGoal. random and longpath are two layouts in Hazard-World-Grid. Easy, Medium and Hard are three levels in SafetyGoal. There are two objects of each hazard type in Easy, four in Medium and six in Hard. Thus, in total, there are 8 hazard objects in Easy, 16 in Medium and 24 in Hard. PPO_Lag refers to the baseline method PPO_Lagrangian, and PPO_CP refers to the proposed method, where the suffix CP means cost prediction. From the results, the proposed method successfully learns a safer policy with the predicted cost. CPO is proposed for robot locomotion tasks. Thus, we do not include it in the first two tasks from Hazard-World-Grid. It is worth noting that in some tasks, the proposed method learns safer policies than the baselines using ground-truth costs. This is because the cost prediction module may mistakenly think of some safe but risky actions as violating the constraints, resulting in a more conservative policy.
@@ 5.1. Experiment Settings
We evaluate the proposed method on two tasks: Hazard-World-Grid Yang et al. (2021) and SafetyGoal Ji et al. (2023) shown in Fig. 2. Each task is accompanied by 20 different constraints to regularize the agents. A full list of detailed constraints for each environment is given in the supplementary material. The pre-trained decoder LM and encoder LM we use are gpt-3.5-turbo Brown et al. (2020) and all-MiniLM-L12-v2 from sentence BERT Reimers and Gurevych (2019b). So, in this section, we also refer to the encoder LM as BERT and the decoder LM as GPT. The detailed constraints, prompts for the decoder LMs to condense constraints and how we transform observation into text-based observations are provided in the supplementary material. On each task, all methods are run for 4 times with different random seeds. The baseline methods we compare here are PPO Schulman et al. (2017), PPO_Lagrangian (PPO_Lag) Ray et al. (2019) and CPO Achiam et al. (2017). PPO only optimizes policies to maximize reward and does not consider the constraints. PPO_Lag enforces constraints with a Lagrangian multiplier to the reward objective and optimizes the policy by solving the equivalent min-max problem. CPO derives policy improvement steps that guarantee both an increase in reward and satisfaction of constraints on costs and is a strong baseline for robot locomotion tasks with constraints. For these two methods, we input the natural language constraints encoded by our encoder LM to the policy network and provide them ground-truth costs to do policy learning, while our proposed method it does not have access to the ground-truth costs at all. Implementation and hyperparameters of the baseline methods can be found here  [^1] https://github.com/PKU-Alignment/Safe-Policy-Optimization. Our method keeps the default hyperparameter and network details of PPO_Lag but use the predicted costs to replace the ground-truth costs.
@@ 5.2. General Results
Figure 4. (a) gives the result of episode reward and (b) gives the result of episode cost as training goes on. The black dashed line in (a) is the converged reward of PPO, and in (b) is the converged cost of ’GPT + BERT + Ground-truth Cost’. The labels are in the format of ’decoder model + encoder model + cost type’. For example, ’GPT + BERT + Predicted Cost’ stands for using GPT as decoder LM, BERT as encoder LM and predicted cost from our cost prediction module for policy learning. (c) gives the cost prediction results when using BERT and LSTM as the encoder model. The cost prediction with LSTM as encoder has very poor performance, which leads to the collapsed training of ’GPT + LSTM + Predicted Cost’ in (a) and (b).
General results on all five tasks are given in Fig. 3. We report episode reward sum and episode cost sum, respectively to show how the methods perform. As we use cost prediction (CP) to replace ground-truth costs, we refer to our method as PPO_CP.
From the results, we can see that compared with PPO that does not consider constraints, PPO_CP can learn safe policies like the baseline methods that require ground-truth costs. This indicates that our cost prediction module is very effective and helpful for learning safe policies. In some tasks, PPO_CP even learns safer policies than the baseline methods, which may be because the cost prediction module mistakenly thinks of some risky but potentially safe actions as violating the constraints (like walking on the edge of the cliff). Thus, the policy is trained to avoid these risky actions, resulting in fewer constraint violations and lower costs. PPO_CP and PPO_Lag achieve even higher rewards than PPO on Random in Hazard-World-Grid. This is because they always tend to find the fastest way to obtain all the goal objects to finish the episode, as they will violate the constraints less if the episode ends fast. And the reward will decay within an episode. Therefore, PPO_CP and PPO_Lag can achieve even higher rewards than PPO.
@@ 5.3.2. Ablation on Encoder LM
BERT is pre-trained on large text corpora so that it can encode natural language input according to their semantic meaning. This is essential for our cost prediction because we need to compare the similarity between the natural language constraint and text-based observation. In this section, we ablate the encoder LM BERT from the proposed method and replace it with a long-short-term-memory (LSTM) network. We remain the other modules of the proposed method, such as GPT as decoder LM, and run an ablation study on layout random in Hazard-World-Grid. As the randomly initialized LSTM cannot tell the difference among constraints, we also pre-train it with the contrastive loss for 10 iterations to make a fair comparison. The ablation results are given in Fig. 4. Labels are in the format of ’decoder model + encoder model + cost type’. For example, ’GPT + BERT + Predicted Cost’ stands for using GPT as decoder LM, BERT as encoder LM and predicted cost from our cost prediction module for policy learning.

### v3

@@ Abstract.
Safe Reinforcement Learning, Natural Language Constraints, Language Models
@@ 1. Introduction
In recent years, great success has been achieved with reinforcement learning (RL) algorithms on multiple domains such as robotic control Kaufmann et al. (2023), resource allocation Zheng et al. (2022) and games Vinyals et al. (2019); Perolat et al. (2022); Du et al. (2019). In real-world settings such as autonomous driving Gu et al. (2022), RL agents are not given complete freedom for safety, fairness or other concerns and must adhere to specific constraints provided by humans. To this end, safe RL is widely studied, following the constrained Markov Decision Process (CMDP) Altman (1999) framework, where agents must satisfy the constraints given by auxiliary costs when optimizing their policies to maximize the objective function.
@@
To learn a safe policy, safe RL algorithms Achiam et al. (2017); Ray et al. (2019); Yang et al. (2020, 2022); Gu et al. (2023) utilize ground-truth cost value given by a well-defined cost function in CMDP. Because the conditions for violating constraints are different, cost functions need to be defined for all constraints case by case and require task-specific domain knowledge, which can be very expensive and inefficient. For example, in a navigation task, the agent should not get close to objects that may cause damage to it. To define such a cost function, one must be aware of all the harmful objects in the environment, otherwise the agent may be harmed by some unexpected hazards. This severely limits safe RL’s application to general and complex tasks where transforming constraints into cost functions can be expensive and requires specific domain knowledge.
@@
One intuitive and easily-accessible approach to regulate safe RL agents is to provide constraints via human language Kaplan et al. (2017); Prakash et al. (2020); Yang et al. (2021); Kazantzidis et al. (2022). By providing natural language constraints, potential end users can easily regularize and interact with agents. However, without a ground-truth cost function, it is difficult to determine whether natural language constraints are violated. To avoid using ground-truth costs, previous method Yang et al. (2021) learns to identify entities relevant to the natural language constraints in the environment so that agents can predict whether the constraints are violated. However, it requires additional computation to model the entities and may result in inaccurate predictions in complex environments. And since the model is trained in a supervised fashion, it is incapable of handling free-form or new natural language constraints from humans.
@@
In this paper, to deal with the problem of constraint violation prediction in safe RL with natural language constraints, we propose to adopt pre-trained language models (LM) Devlin et al. (2018); Brown et al. (2020) to comprehend and encode the constraints and predict costs for policy learning. Specifically, a decoder-based LM (such as GPT Brown et al. (2020)) is adopted to eliminate ambiguity and redundancy and extract semantic meaning from the constraints. As decoder LMs are pre-trained on a very large corpus of data and align with human values, they are able to generate high-quality constraints with condensed semantic meaning if we prompt them properly. To condition the policy which is parameterized by neural networks with natural language constraints, an encoder-based LM (such as BERT Devlin et al. (2018)) encodes constraints into embeddings according to their semantic meaning. We also use the encoder LM to encode the text-based observations. By computing the semantic similarities between constraint embeddings and description embeddings, we can determine whether the constraint is violated. To distinct constraint embeddings from each other, we adopt a contrastive loss to fine-tune the encoder LM so that constraints have similar embeddings only when they have similar semantic meaning.
@@
    We introduce pre-trained LMs to safe RL with natural language constraints to replace the recurrent neural network in previous works, which can only handle structured or fixed-form natural language constraints.
@@ 4.1. Cost Prediction Module
We give the structure of the cost prediction module in Fig. 1. In general, to predict whether the constraint is violated, we adopt a decoder LM (GPT) to eliminate ambiguity and condense semantic meaning in free-form natural language constraint $c$ and an encoder LM (BERT) to extract semantic information from the condensed natural language constraint $\overline{c}$ and transform the constraints into embedding $h_{c}$. Condensed constraint $\overline{c}$ aligns with the original constraint $c$ on which entity or behaviour the agents are prohibited from. This is necessary as free-form constraint $c$ from humans may be verbose or semantically vague, which will severely affect the cost prediction result. Since decoder LM GPT aligns with human values Ziegler et al. (2019), it can eliminate ambiguity within the human natural language constraints and better summarize the constraints. We empirically show in the experiment section that without the decoder LM, the performance of the cost prediction module will drop significantly. The prompt we use for the decoder LM is provided in the supplementary material.
@@ 4.2. Policy Training
Algorithm 1 PPO with LM-based Cost Prediction (PPO-CP)
@@
$L=L^{CLIP}-\alpha L^{CLIP}_{C}-c_{1}L^{VF}-c_{2}L^{CVF},$ (8)
@@ 5.1. Experiment Settings
Figure 3. Experiment results on Hazard-World-Grid and SafetyGoal. random and longpath are two layouts in Hazard-World-Grid. Easy, Medium and Hard are three levels in SafetyGoal. There are two objects of each hazard type in Easy, four in Medium and six in Hard. Thus, in total, there are 8 hazard objects in Easy, 16 in Medium and 24 in Hard. PPO-Lag refers to the baseline method PPO-Lagrangian, and PPO-CP refers to the proposed method, where the suffix CP means cost prediction. From the results, the proposed method successfully learns a safer policy with the predicted cost. CPO is proposed for robot locomotion tasks. Thus, we do not include it in the first two tasks from Hazard-World-Grid. It is worth noting that in some tasks, the proposed method learns safer policies than the baselines using ground-truth costs. This is because the cost prediction module may mistakenly think of some safe but risky actions as violating the constraints, resulting in a more conservative policy.
We evaluate the proposed method on two tasks: Hazard-World-Grid Yang et al. (2021) and SafetyGoal Ji et al. (2023) shown in Fig. 2. Each task is accompanied by 20 different constraints to regularize the agents. A full list of detailed constraints for each environment is given in the supplementary material. The pre-trained decoder LM and encoder LM we use are gpt-3.5-turbo Brown et al. (2020) and all-MiniLM-L12-v2 from sentence BERT Reimers and Gurevych (2019b). So, in this section, we also refer to the encoder LM as BERT and the decoder LM as GPT. The detailed constraints, prompts for the decoder LMs to condense constraints and how we transform observation into text-based observations are provided in the supplementary material. On each task, all methods are run for 4 times with different random seeds. The baseline methods we compare here are PPO Schulman et al. (2017), PPO-Lagrangian (PPO-Lag) Ray et al. (2019) and CPO Achiam et al. (2017). PPO only optimizes policies to maximize reward and does not consider the constraints. PPO-Lag enforces constraints with a Lagrangian multiplier to the reward objective and optimizes the policy by solving the equivalent min-max problem. CPO derives policy improvement steps that guarantee both an increase in reward and satisfaction of constraints on costs and is a strong baseline for robot locomotion tasks with constraints. For these two methods, we input the natural language constraints encoded by our encoder LM to the policy network and provide them ground-truth costs to do policy learning, while our proposed method it does not have access to the ground-truth costs at all. Implementation and hyperparameters of the baseline methods can be found here  [^1] https://github.com/PKU-Alignment/Safe-Policy-Optimization. Our method keeps the default hyperparameter and network details of PPO-Lag but use the predicted costs to replace the ground-truth costs.
@@ 5.2. General Results
Figure 4. (a) gives the result of episode reward and (b) gives the result of episode cost as training goes on. The labels are in the format of ’decoder model + encoder model + cost type’. For example, ’GPT + BERT + Predicted Cost’ stands for using GPT as decoder LM, BERT as encoder LM and predicted cost from our cost prediction module for policy learning. ’GPT + BERT + Predicted Cost’ is our proposed method PPO-CP, and ’GPT + BERT + Ground-truth Cost’ is PPO-Lag in Fig. 3. (c) gives the cost prediction results when using BERT and LSTM as the encoder model. The cost prediction with LSTM as encoder has very poor performance, which leads to the collapsed training of ’GPT + LSTM + Predicted Cost’ in (a) and (b).
General results on all five tasks are given in Fig. 3. We report episode reward sum and episode cost sum, respectively to show how the methods perform. As we use cost prediction (CP) to replace ground-truth costs, we refer to our method as PPO-CP.
From the results, we can see that compared with PPO that does not consider constraints, PPO-CP can learn safe policies like the baseline methods that require ground-truth costs. This indicates that our cost prediction module is very effective and helpful for learning safe policies. In some tasks, PPO-CP even learns safer policies than the baseline methods, which may be because the cost prediction module mistakenly thinks of some risky but potentially safe actions as violating the constraints (like walking on the edge of the cliff). Thus, the policy is trained to avoid these risky actions, resulting in fewer constraint violations and lower costs. PPO-CP and PPO-Lag achieve even higher rewards than PPO on Random in Hazard-World-Grid. This is because they always tend to find the fastest way to obtain all the goal objects to finish the episode, as they will violate the constraints less if the episode ends fast. And the reward will decay within an episode. Therefore, PPO-CP and PPO-Lag can achieve even higher rewards than PPO.
@@ 5.3.2. Ablation on Encoder LM
BERT is pre-trained on large text corpora so that it can encode natural language input according to their semantic meaning. This is essential for our cost prediction because we need to compare the similarity between the natural language constraint and text-based observation. In this section, we ablate the encoder LM BERT from the proposed method and replace it with a long-short-term-memory (LSTM) network. We remain the other modules of the proposed method, such as GPT as decoder LM, and run an ablation study on layout random in Hazard-World-Grid. As the randomly initialized LSTM cannot tell the difference among constraints, we also pre-train it with the contrastive loss for 10 iterations to make a fair comparison. The ablation results are given in Fig. 4. Labels are in the format of ’decoder model + encoder model + cost type’. For example, ’GPT + BERT + Predicted Cost’ stands for using GPT as decoder LM, BERT as encoder LM and predicted cost from our cost prediction module for policy learning. The decoder LM extracts semantic meaning from the constraints and is adopted in all methods except PPO, because it ignores the constraints. It is worth noting although PPO-Lag has access to the ground-truth cost, it still needs an Encoder LM to encode the natural language constraints, so that the constraints can be put into and constrain the policy network.